It's been an embarrassingly long time since my undergraduate statistics class, so the math might not be 100% correct, but I'd like to attempt it anyway, be sure to tell me if I'm wrong:
Worst case, you fly twice a month for 20 years, that's 480 flights. Each flight is a round trip, so baggage is entered into the airlines systems 960 times in 20 years.
If the probability that an airline gets your baggage to you each time is 99.99%
.9999^960 = .9085
1-.9085 = .0915
So, if my calculations are correct, in 20 years time, there's a 9.15% of chance your bags having be lost at least once. Assuming that the odds of an airline losing all bags are the same - and they are obviously not.